OpenStudy (anonymous):

What is twelve twenty-fifths divided by four-fifths? one and two-thirds one and three-fifths three-fifths three twenty-fifths

OpenStudy (anonymous):

About tree fitty

OpenStudy (anonymous):

um thx

OpenStudy (anonymous):

is it three fiths

OpenStudy (anonymous):

if your gonna joke get the heck off my question

OpenStudy (anonymous):

zabali plz help

OpenStudy (anonymous):

stop messing around i take math seriously you green fat ugly ogar

OpenStudy (anonymous):

whats the answer im srry

OpenStudy (anonymous):

About Tree Fitty. <3

OpenStudy (anonymous):

thx god zambali is helping me i know he takes math seriously

OpenStudy (anonymous):

Inb4 Tree fitty. Cmon bro

OpenStudy (anonymous):

12/25 divided 4/5 =\[\frac{ 12 }{ 25 }\div \frac{ 4 }{ 5 }=\frac{ 12 }{ 25 }\times \frac{ 5 }{ 4 }=\frac{ 60 }{ 100 }=\frac{ 3 }{ 5 }\] so it is three fifth..

OpenStudy (anonymous):

just clik best response button on my coment..

OpenStudy (anonymous):

i did thx

OpenStudy (anonymous):

you welocme..